# Regularized Q-Learning with Linear Function Approximation

### v1

@@ 9 Proofs of Theorems
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left\|\omega^{*}(\theta^{t})-\omega^{t+1}\right\|^{2}=\mathcal{O}\left(T^{-1/2}\right).$
See Appendix 10.3 for a detailed proof. As a result, by Jensen’s inequality, we can derive
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left\|\omega^{*}(\theta^{t})-\omega^{t+1}\right\|\leq$ $\sqrt{\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left\|\omega^{*}(\theta^{t})-\omega^{t+1}\right\|^{2}}$
$=$ $\mathcal{O}\left(T^{-1/4}\right).$ (20)
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\left\|\bar{\nabla}f\left(\theta^{t},\omega^{t+1}\right)\right\|=\mathcal{O}\left(T^{-1/4}\right).$
See Appendix 10.4 for a detailed proof.
Finally, it is straightforward to derive (13). By the triangle inequality and Lemma 8.9, we have
$\frac{1}{T}$ $\sum_{t=0}^{T-1}\left\|{\nabla}J\left(\theta^{t}\right)\right\|$
$\leq\frac{1}{T}\sum_{t=0}^{T-1}\left(\left\|{\nabla}J\left(\theta^{t}\right)-\bar{\nabla}f\left(\theta^{t},\omega^{t+1}\right)\right\|+\left\|\bar{\nabla}f\left(\theta^{t},\omega^{t+1}\right)\right\|\right)$
$\leq\frac{2}{T}\sum_{t=0}^{T-1}\left\|\omega^{*}\left(\theta^{t}\right)-\omega^{t+1}\right\|+\frac{1}{T}\sum_{t=0}^{T-1}\left\|\bar{\nabla}f\left(\theta^{t},\omega^{t+1}\right)\right\|$
$=\mathcal{O}\left(T^{-1/4}\right).$
@@
For the first term in (9.2), due to regularized Bellman equation and the $\gamma$-contraction property of the smooth truncated optimal regularized Bellman operator $\mathcal{B}_{\tau,\delta}$ in $\ell_{\infty}$-norm, we obtain
@@
where the last inequality follows from the fact $\|V_{\pi,\tau}\|_{\infty}\leq\frac{R_{\max}+\tau B}{1-\gamma}=\delta_{0}$ for any policy $\pi$ and (2). Next, we can bound the second term in (9.2) by the definition of approximation error in (14):
@@
Finally, we bound the last term in (9.2). Under Assumption 3.4 and given the expression of $\nabla J(\cdot)$, we have
$\left\|\theta^{t}-\omega^{*}\left(\theta\right)\right\|\leq\frac{1}{\sigma_{\min}}\left\|\nabla J\left(\theta^{t}\right)\right\|$
@@
Substituting (9.2), (23) and (9.2) into (9.2) leads to
@@
Taking full expectation and summing from $t=0$ to $t=T-1$ for the above inequality completes the proof of (5.7).
For the second bound in Theorem 5.7, given the definition of $V_{\tau}^{t}$, we have
@@
For the first term in (9.2), we utilize the properties of the convex conjugate $G^{*}$ of the regularizer:
@@
For the second term in (9.2), the bound can be developed similarly. Given definition of $\pi_{\theta}$ in (4), for $s\in\mathcal{S}$, we have
@@
Here, our focus is on providing a bound for the first term, as bounds for the second and third terms have already been established in (23) and (9.2). Given the regularized Bellman equation, we can derive
@@
where the last inequality follows from the $1$-Lipschitz property of $\mathcal{K}_{\delta}$. Substituting (9.2), (23) and (9.2) into (9.2) leads to
@@
Combining the above inequality with (9.2) and using the result in (5.7) complete the proof of Theorem 5.7.
10 Proofs of Other Lemmas
10.1 Proof of Lemma 5.1
@@ 10 Proofs of Other Lemmas
where the third inequality is due to Lemma H1 and Lemma 8.5, and the last inequality follows from the definition of $\hat{\Sigma}_{\tau,\delta,\theta}$ in (3.2).
@@
10.2 Proof of Lemma 5.2
By the definition of $h_{g}^{t}$ in (10a), we have
@@
10.3 Proof of Lemma 9.1
For $t\geq T_{\alpha}$, following the update of $\omega^{t+1}$ in (11a) and taking the conditional expectation given filtration $\mathcal{F}_{t}$ yields
$\mathbb{E}$ $\left[\left\|\omega^{*}(\theta^{t})-\omega^{t+1}\right\|^{2}\mid\mathcal{F}_{t}\right]$
$\leq\mathbb{E}\left[\left\|\omega^{*}\left(\theta^{t}\right)-\omega^{t}+\beta h_{g}^{t}\right\|^{2}\mid\mathcal{F}_{t}\right]$
$=\left\|\omega^{*}\left(\theta^{t}\right)-\omega^{t}\right\|^{2}$
$\quad+\mathbb{E}\Big{[}2\beta\left\langle\omega^{*}\left(\theta^{t}\right)-\omega^{t},\tilde{h}_{g}^{t}-\tilde{h}_{g}^{t}+h_{g}^{t}\right\rangle+\beta^{2}\left\|h_{g}^{t}\right\|^{2}\mid\mathcal{F}_{t}\Big{]}$
@@
$=\left\|\omega^{*}\left(\theta^{t}\right)-\omega^{t}\right\|^{2}+2\beta\left\langle\omega^{*}\left(\theta^{t}\right)-\omega^{t},\nabla_{\omega}g(\theta^{t},\omega^{t})\right\rangle$
$\quad+2\beta\left\langle\omega^{*}\left(\theta^{t}\right)-\omega^{t},\mathbb{E}\left[h_{g}^{t}-\tilde{h}_{g}^{t}\mid\mathcal{F}_{k}\right]\right\rangle$
$\quad\quad+\beta^{2}\mathbb{E}\left[\left\|h_{g}^{t}\right\|^{2}\mid\mathcal{F}_{t}\right]$
$\leq 2\beta\left(g(\theta^{t},\omega^{*}\left(\theta^{t}\right))-g(\theta^{t},\omega^{t})\right)$
$\ \ \ \ \ \ +(1-\beta\lambda_{g})\left\|\omega^{*}\left(\theta^{t}\right)-\omega^{t}\right\|^{2}$
$\quad\quad\quad+2\beta\left\langle\omega^{*}\left(\theta^{t}\right)-\omega^{t},\mathbb{E}\left[h_{g}^{t}-\tilde{h}_{g}^{t}\mid\mathcal{F}_{k}\right]\right\rangle$
$\quad\quad\quad\quad+\beta^{2}\mathbb{E}\left[\left\|h_{g}^{t}\right\|^{2}\mid\mathcal{F}_{t}\right]$
$\leq(1-\beta\lambda_{g})\|\omega^{*}\left(\theta^{t}\right)-\omega^{t}\|^{2}$
$\quad+2\beta\left\langle\omega^{*}\left(\theta^{t}\right)-\omega^{t},\mathbb{E}\left[h_{g}^{t}-\tilde{h}_{g}^{t}\mid\mathcal{F}_{k}\right]\right\rangle$
$\quad\quad+\beta^{2}\mathbb{E}\left[\left\|h_{g}^{t}\right\|^{2}\mid\mathcal{F}_{t}\right],$ (29)
@@
where the first inequality follows from the property of the projection $\mathcal{P}$ and the third inequality is due to Proposition 3.3.
@@
We first analyze the first term in (10.3). If $\beta\geq 1/\lambda_{g}$, this term can be upper bounded by 00 and the subsequent results are easy to derive. If otherwise, meaning $\beta<1/\lambda_{g}$, we have
$(1-\beta\lambda_{g})\|\omega^{*}\left(\theta^{t}\right)-\omega^{t}\|^{2}$
$=\left(1-\beta\lambda_{g}\right)\left\|\omega^{*}\left(\theta^{t}\right)-\omega^{*}(\theta^{t-1})+\omega^{*}(\theta^{t-1})-\omega^{t}\right\|^{2}$
$\leq\left(1-\beta\lambda_{g}\right)\left(1+c\right)\left\|\omega^{*}\left(\theta^{t}\right)-\omega^{*}(\theta^{t-1})\right\|^{2}$
$\ \ \ \ \ \ +\left(1-\beta\lambda_{g}\right)\left(1+\frac{1}{c}\right)\left\|\omega^{*}(\theta^{t-1})-\omega^{t}\right\|^{2}$
$\leq\left(1-\beta\lambda_{g}\right)\left((1+c)\frac{4\alpha^{2}}{\lambda_{g}^{2}}+\left(1+\frac{1}{c}\right)\left\|\omega^{*}(\theta^{t-1})-\omega^{t}\right\|^{2}\right)$
$\leq\frac{8\alpha^{2}}{\beta\lambda_{g}^{3}}+\left(1-\frac{\beta\lambda_{g}}{2}\right)\left\|\omega^{*}(\theta^{t-1})-\omega^{t}\right\|^{2},$ (30)
where the first inequality is due to Young’s inequality (and holds for any $c>0$), the second one follows from combining Lemma H1 and Lemma 8.7, and the last one is obtained by setting $c=\frac{2(1-\beta\lambda_{g})}{\beta\lambda_{g}}$.
For the second term in (10.3), we can bound it as:
$\left\langle\omega^{*}\left(\theta^{t}\right)-\omega^{t},\mathbb{E}\left[h_{g}^{t}-\tilde{h}_{g}^{t}\mid\mathcal{F}_{k}\right]\right\rangle$
$\quad\leq\left\|\omega^{*}\left(\theta^{t}\right)-\omega^{t}\right\|\cdot\left\|\mathbb{E}\left[h_{g}^{t}-\tilde{h}_{g}^{t}\mid\mathcal{F}_{k}\right]\right\|$
$\quad\leq\frac{2\left(R_{\max}+\delta\right)}{\lambda_{g}}\left\|\mathbb{E}\left[h_{g}^{t}-\tilde{h}_{g}^{t}\mid\mathcal{F}_{k}\right]\right\|$
$\quad\leq\frac{4\left(R_{\max}+\delta\right)^{2}}{\lambda_{g}^{2}}$
×$\quad\quad\times d_{TV}\left(\mathbb{P}\left(\left(s_{t},a_{t},s_{t}^{\prime}\right)\in\cdot\mid s_{0}\sim\mu_{0},\pi_{\mathrm{bhv}},P\right),\mathcal{D}\right)$
$\quad\leq\frac{4\alpha\left(R_{\max}+\delta\right)^{2}}{\lambda_{g}^{2}},$ (31)
where the third inequality follows the definition of the total-variation norm and Lemma 5.2, and the last inequality uses Lemma 1 in [48].
The last term in (10.3) can be bounded by Lemma 5.2 as
$\mathbb{E}\left[\left\|h_{g}^{t}\right\|^{2}\mid\mathcal{F}_{k}\right]\leq\frac{4\left(R_{\max}+\delta\right)^{2}}{\lambda_{g}^{2}}.$ (32)
Substituting (10.3), (10.3) and (32) into (10.3) yields
$\mathbb{E}$ $\left[\left\|\omega^{*}(\theta^{t})-\omega^{t+1}\right\|^{2}\mid\mathcal{F}_{t}\right]$
$\quad\quad\leq\left(1-\frac{\beta\lambda_{g}}{2}\right)\left\|\omega^{*}(\theta^{t-1})-\omega^{t}\right\|^{2}+\frac{8\alpha^{2}}{\beta\lambda_{g}^{3}}$
$\quad\quad\quad\quad+\frac{\left(8\alpha\beta+4\beta^{2}\right)\left(R_{\max}+\delta\right)^{2}}{\lambda_{g}^{2}}.$ (33)
Taking the full expectation of the above inequality and summing it from $t=T_{\alpha}+1$ to $t=T$ gives
$\sum_{t=T_{\alpha}+1}^{T}\mathbb{E}$ $\left\|\omega^{*}(\theta^{t})-\omega^{t+1}\right\|^{2}$
@@
$\leq$ $\left(1-\frac{\beta\lambda_{g}}{2}\right)\sum_{t=T_{\alpha}+1}^{T}\mathbb{E}\left\|\omega^{*}(\theta^{t-1})-\omega^{t}\right\|^{2}+\frac{8T\alpha^{2}}{\beta\lambda_{g}^{3}}$
$\quad+\frac{\left(8T\alpha\beta+4T\beta^{2}\right)\left(R_{\max}+\delta\right)^{2}}{\lambda_{g}^{2}}.$
Rearranging the above equality leads to
$\frac{1}{T}\sum_{t=T_{\alpha}}^{T-1}\mathbb{E}$ $\left\|\omega^{*}(\theta^{t})-\omega^{t+1}\right\|^{2}$
@@
$\leq\frac{2}{\beta\lambda_{g}T}\mathbb{E}\left\|\omega^{*}(\theta^{T_{\alpha}})-\omega^{T_{\alpha}+1}\right\|^{2}+\frac{16\alpha^{2}}{\beta^{2}\lambda_{g}^{4}}$
$\qquad+\frac{\left(16\alpha+8\beta\right)\left(R_{\max}+\delta\right)^{2}}{\lambda_{g}^{3}}$
@@
$\leq\frac{8\left(R_{\max}+\delta\right)^{2}}{T\beta\lambda_{g}^{3}}+\frac{16\alpha^{2}}{\beta^{2}\lambda_{g}^{4}}$
@@
$\qquad+\frac{\left(16\alpha+8\beta\right)\left(R_{\max}+\delta\right)^{2}}{\lambda_{g}^{3}}.$ (34)

### v3

@@ Appendix B Proofs of Theorems
We define $T_{\beta}$ in a similar way:
$T_{\beta}\triangleq\min\left\{t\in\mathbb{N}^{+}\mid\kappa\rho^{t}\leq\beta\right\}.$ (20)
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left\|\omega^{*}(\theta^{t})-\omega^{t+1}\right\|^{2}=\mathcal{O}\left(T^{-1/2}\log T\right).$
See Appendix C.4 for a detailed proof. As a result, by Jensen’s inequality, we can derive
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left\|\omega^{*}(\theta^{t})-\omega^{t+1}\right\|\leq\sqrt{\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left\|\omega^{*}(\theta^{t})-\omega^{t+1}\right\|^{2}}=\mathcal{O}\left(T^{-1/4}(\log T)^{1/2}\right).$ (21)
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\left\|\bar{\nabla}_{\theta}f\left(\theta^{t},\omega^{t+1}\right)\right\|=\mathcal{O}\left(T^{-1/4}(\log T)^{1/2}\right).$
See Appendix C.5 for a detailed proof.
Finally, it is straightforward to derive (14). By the triangle inequality and Lemma A5, we have
$\frac{1}{T}\sum_{t=0}^{T-1}\left\|{\nabla}J\left(\theta^{t}\right)\right\|$ $\leq\frac{1}{T}\sum_{t=0}^{T-1}\left(\left\|{\nabla}J\left(\theta^{t}\right)-\bar{\nabla}_{\theta}f\left(\theta^{t},\omega^{t+1}\right)\right\|+\left\|\bar{\nabla}_{\theta}f\left(\theta^{t},\omega^{t+1}\right)\right\|\right)$
$\leq\frac{2}{T}\sum_{t=0}^{T-1}\left\|\omega^{*}\left(\theta^{t}\right)-\omega^{t+1}\right\|+\frac{1}{T}\sum_{t=0}^{T-1}\left\|\bar{\nabla}_{\theta}f\left(\theta^{t},\omega^{t+1}\right)\right\|$
$=\mathcal{O}\left(T^{-1/4}(\log T)^{1/2}\right).$
@@
For the first term in (B.2), due to regularized Bellman equation and the $\gamma$-contraction property of the smooth truncated optimal regularized Bellman operator $\mathcal{B}_{\tau,\delta}$ in $\ell_{\infty}$-norm, we obtain
@@
where the last inequality follows from the fact $\|V_{\pi,\tau}\|_{\infty}\leq\frac{R_{\max}+\tau B}{1-\gamma}=\delta_{0}$ for any policy $\pi$ and (4). Next, we can bound the second term in (B.2) by the definition of approximation error in (15):
@@
Finally, we bound the last term in (B.2). Under Assumption 3.2 and given the expression of $\nabla J(\cdot)$, we have
$\left\|\theta^{t}-\omega^{*}\left(\theta^{t}\right)\right\|\leq\frac{1}{\sigma_{\min}}\left\|\nabla J\left(\theta^{t}\right)\right\|$
@@
Substituting (B.2), (24) and (B.2) into (B.2) leads to
@@
Taking full expectation and summing from $t=0$ to $t=T-1$ for the above inequality completes the proof of (5.7). For the second bound in Theorem 5.7, given the definition of $V_{\tau}^{t}$, we have
@@
For the first term in (B.2), we utilize the properties of the convex conjugate $G^{*}$ of the regularizer:
@@
For the second term in (B.2), the bound can be developed similarly. Given definition of $\pi_{\theta}$ in (3), for $s\in\mathcal{S}$, we have
@@
Here, our focus is on providing a bound for the first term, as bounds for the second and third terms have already been established in (24) and (B.2). Given the regularized Bellman equation, we can derive
@@
where the last inequality follows from the $1$-Lipschitz property of $\mathcal{K}_{\delta}$. Substituting (B.2), (24) and (B.2) into (B.2) leads to
@@
Combining the above inequality with (B.2) and using the result in (5.7) complete the proof of Theorem 5.7.
Appendix C Proofs of Other Lemmas
C.1 Proof of Lemma 4.1
With $\delta=\delta_{0}$, we now show $\sup_{t}\max_{a\in\mathcal{A}}\big{\{}|\phi(s_{t+1},a)^{\top}\theta^{t}|\big{\}}<\infty$ w.p. 1. By contradiction, we consider two cases separately:
  • a.
    $\max_{a\in\mathcal{A}}\big{\{}|\phi(s_{t+1},a)^{\top}\theta^{t}|\big{\}}>\delta_{0}+\tau B$ eventually.
  • b.
    $\max_{a\in\mathcal{A}}\big{\{}|\phi(s_{t+1},a)^{\top}\theta^{t}|\big{\}}<\delta_{0}+\tau B$ infinitely often.
Let us consider case (a), i.e. there is a $T<\infty$ such that $\max_{a\in\mathcal{A}}\big{\{}|\phi(s_{t+1},a)^{\top}\theta^{t}|\big{\}}>\delta_{0}+\tau B$ for all $t\geq T$. By Proposition 2.1.(i), we have $|G^{*}_{\tau}(\hat{Q}_{\theta^{t}}(s_{t+1},\cdot))|\geq\delta_{0}$, hence, truncation is implemented for all $t\geq T$ and:
$h_{g}^{t}$ $=\phi(s_{t},a_{t})\left(\phi(s_{t},a_{t})^{\top}\omega^{t}-R(s_{t},a_{t})-\gamma\delta_{0}\right)$
for all $t\geq T$. By ergodicity assumption, $\omega^{t}\rightarrow\omega^{*}$ and $\theta^{t}\rightarrow\omega^{*}$ where
$\omega^{*}$ $\triangleq\Sigma^{-1}\mathbb{E}_{\mathcal{D}}\left[\phi(s,a)\Big{(}R(s,a)+\gamma\delta_{0}\Big{)}\right]$
However, $|\hat{G}_{\tau,\delta}\big{(}\hat{Q}_{\omega^{*}}(s_{t+1},\cdot))\big{)}|<\delta_{0}$. A contradiction. For case (b), let $\{\theta^{t(n)}\}$ be a subsequence such that $\max_{a\in\mathcal{A}}\big{\{}|\phi(s_{t+1},a)^{\top}\theta^{t(n)}|\big{\}}\rightarrow\infty$. Let $\tau(n)$ be defined as
$\tau(n)$ $\triangleq\max\big{\{}t\leq t(n):\max_{a\in\mathcal{A}}|\phi(s_{t+1},a)^{\top}\theta^{t}|\leq\delta_{0}+\tau B\big{\}}$
By Lemma A.4 in the paper, for all $t>0$, $\|\theta^{t+1}-\theta^{t}\|\leq 2\alpha$. Hence,
$t(n)-\tau(n)\geq\frac{\max_{a\in\mathcal{A}}|\phi(s_{t+1},a)^{\top}\theta^{t}|-(\delta_{0}+\tau B)}{2\alpha}$
By hypothesis $\max_{a\in\mathcal{A}}\big{\{}|\phi(s_{t+1},a)^{\top}\theta^{t(n)}|\big{\}}\rightarrow\infty$, hence $t(n)-\tau(n)\rightarrow\infty$. As in the proof of case (a),
$h_{g}^{t}$ $=\phi(s_{t},a_{t})\left(\phi(s_{t},a_{t})^{\top}\omega^{t}-R(s_{t},a_{t})-\gamma\delta_{0}\right)~{}~{}~{}~{}~{}~{}t\in(\tau(n),t(n)]$
and $\lim_{n}\omega^{t(n)}=\omega^{*}$ and $\lim_{n}\theta^{t(n)}=\omega^{*}$. A contradiction.
C.2 Proof of Lemma 5.1
@@ Appendix C Proofs of Other Lemmas
where the third inequality is due to Lemma A1 and Lemma A3, and the last inequality follows from the definition of $\hat{\Sigma}_{\tau,\delta,\theta}$ in (3.3).
@@
C.3 Proof of Lemma 5.2
By the definition of $h_{g}^{t}$ in (10), we have
@@
C.4 Proof of Lemma B1
To start, we define
$o^{t}_{g}(\theta,\omega)\triangleq\phi(s_{t},a_{t})(\phi(s_{t},a_{t})^{\top}\omega-R(s_{t},a_{t})-\gamma\hat{G}_{\tau,\delta}(\hat{Q}_{\theta}(s^{\prime}_{t},\cdot))),$
and
$\xi^{t}_{g}(\theta,\omega)\triangleq\langle\omega^{*}(\theta)-\omega,o^{t}_{g}(\theta,\omega)-\nabla_{\omega}g(\theta,\omega)\rangle.$
$\xi^{t}_{g}$ has the following properties.
Lemma C1.
For $t\geq 0$, $\xi^{t}_{g}(\theta,\omega)$ is
  1. 1.
    $\frac{8(R_{\max}+\delta)}{\lambda_{g}}$-Lipchitz in $\omega$ for $\omega\in\Omega$,
  2. 2.
    and $\frac{4(R_{\max}+\delta)(1+\lambda_{g})}{\lambda_{g}^{2}}$-Lipschitz in $\theta$.
The proof of Lemma C1 is in Appendix C.6.
For $t\geq T_{\alpha}$, following the update of $\omega^{t+1}$ in (12a) and taking the expectation yields
$\mathbb{E}\left[\left\|\omega^{*}(\theta^{t})-\omega^{t+1}\right\|^{2}\right]$
$\leq$ $\mathbb{E}\left[\left\|\omega^{*}\left(\theta^{t}\right)-\omega^{t}+\beta h_{g}^{t}\right\|^{2}\right]$
@@ C.4 Proof of Lemma B1
$=$ $\mathbb{E}\left[\left\|\omega^{*}\left(\theta^{t}\right)-\omega^{t}\right\|^{2}\right]+2\beta\mathbb{E}\left[\left\langle\omega^{*}\left(\theta^{t}\right)-\omega^{t},\nabla_{\omega}g(\theta^{t},\omega^{t})\right\rangle\right]$
$+2\beta\mathbb{E}\left[\xi^{t}_{g}(\theta^{t},\omega^{t})\right]+\beta^{2}\mathbb{E}\left[\left\|h_{g}^{t}\right\|^{2}\right]$
$\leq$ $(1-\beta\lambda_{g})\mathbb{E}\left[\left\|\omega^{*}\left(\theta^{t}\right)-\omega^{t}\right\|^{2}\right]+2\beta\mathbb{E}\left[\xi^{t}_{g}(\theta^{t},\omega^{t})\right]$
$+\beta^{2}\mathbb{E}\left[\left\|h_{g}^{t}\right\|^{2}\right],$ (30)
@@
where the first inequality follows from the property of the projection $\mathcal{P}$ and the last inequality is due to Proposition E.1.
@@
We first analyze the first term in (C.4). If $\beta\geq 1/\lambda_{g}$, this term can be upper bounded by 00 and the subsequent results are easy to derive. If otherwise, meaning $\beta<1/\lambda_{g}$, we have
$(1-\beta\lambda_{g})\mathbb{E}\left[\|\omega^{*}\left(\theta^{t}\right)-\omega^{t}\|^{2}\right]$
$=\left(1-\beta\lambda_{g}\right)\mathbb{E}\left[\left\|\omega^{*}\left(\theta^{t}\right)-\omega^{*}(\theta^{t-1})+\omega^{*}(\theta^{t-1})-\omega^{t}\right\|^{2}\right]$
$\leq\left(1-\beta\lambda_{g}\right)\left(1+c\right)\mathbb{E}\left[\left\|\omega^{*}\left(\theta^{t}\right)-\omega^{*}(\theta^{t-1})\right\|^{2}\right]$
@@
$\ \ \ \ \ \ +\left(1-\beta\lambda_{g}\right)\left(1+\frac{1}{c}\right)\mathbb{E}\left[\left\|\omega^{*}(\theta^{t-1})-\omega^{t}\right\|^{2}\right]$
@@
$\leq\left(1-\beta\lambda_{g}\right)\left((1+c)\frac{4\alpha^{2}}{\lambda_{g}^{2}}+\left(1+\frac{1}{c}\right)\mathbb{E}\left[\left\|\omega^{*}(\theta^{t-1})-\omega^{t}\right\|^{2}\right]\right)$
@@
$\leq\frac{8\alpha^{2}}{\beta\lambda_{g}^{3}}+\left(1-\frac{\beta\lambda_{g}}{2}\right)\mathbb{E}\left[\left\|\omega^{*}(\theta^{t-1})-\omega^{t}\right\|^{2}\right],$ (31)
where the first inequality is due to Young’s inequality (and holds for any $c>0$), the second one follows from combining Lemma A1 and Lemma A4, and the last one is obtained by setting $c=\frac{2(1-\beta\lambda_{g})}{\beta\lambda_{g}}$.
For the second term in (C.4), we first define a random variable $\tilde{\xi}_{g}(\theta,\omega)$:
$\tilde{\xi}_{g}(\theta,\omega)\triangleq\langle\omega^{*}(\theta)-\omega,\phi(\tilde{s},\tilde{a})(\phi(\tilde{s},\tilde{a})^{\top}\omega-R(\tilde{s},\tilde{a})-\gamma\hat{G}_{\tau,\delta}(\hat{Q}_{\theta}(\tilde{s}^{\prime},\cdot)))-\nabla_{\omega}g(\theta,\omega)\rangle,$
where $(\tilde{s},\tilde{a},\tilde{s}^{\prime})\sim\mathcal{D}$. Given $t\geq T_{\alpha}$, we have
$\mathbb{E}\left[\xi^{t}_{g}(\theta^{t},\omega^{t})\mid\mathcal{F}_{t-T_{\alpha}}\right]=$ $\mathbb{E}\left[\xi^{t}_{g}(\theta^{t},\omega^{t})-\xi^{t}_{g}(\theta^{t},\omega^{t-T_{\alpha}})\mid\mathcal{F}_{t-T_{\alpha}}\right]$
$+\mathbb{E}\left[\xi^{t}_{g}(\theta^{t},\omega^{t-T_{\alpha}})-\xi^{t}_{g}(\theta^{t-T_{\alpha}},\omega^{t-T_{\alpha}})\mid\mathcal{F}_{t-T_{\alpha}}\right]$
$+\mathbb{E}\left[\xi^{t}_{g}(\theta^{t-T_{\alpha}},\omega^{t-T_{\alpha}})-\tilde{\xi}_{g}(\theta^{t-T_{\alpha}},\omega^{t-T_{\alpha}})\mid\mathcal{F}_{t-T_{\alpha}}\right]$
@@
$\leq$ $\frac{8(R_{\max}+\delta)}{\lambda_{g}}\mathbb{E}\left[\left\|\omega^{t}-\omega^{t-T_{\alpha}}\right\|\mid\mathcal{F}_{t-T_{\alpha}}\right]$
$+\frac{4(R_{\max}+\delta)(1+\lambda_{g})}{\lambda_{g}^{2}}\mathbb{E}\left[\left\|\theta^{t}-\theta^{t-T_{\alpha}}\right\|\mid\mathcal{F}_{t-T_{\alpha}}\right]$
$+\frac{4\left(R_{\max}+\delta\right)^{2}}{\lambda_{g}^{2}}d_{TV}\left(\mathbb{P}\left(\left(s_{t},a_{t},s_{t}^{\prime}\right)\in\cdot\mid s_{t-T_{\alpha}},\pi_{\mathrm{bhv}},P\right),\mathcal{D}\right)$
$\leq$ $\frac{(16\beta T_{\alpha}+4\alpha)(R_{\max}+\delta)^{2}+8\alpha T_{\alpha}(R_{\max}+\delta)(1+\lambda_{g})}{\lambda_{g}^{2}},$ (32)
where the first inequality is due to Lemma C1, Assumption 2.2, and Lemma 1 in [35].
The last term in (C.4) can be bounded by Lemma 5.2 as
$\mathbb{E}\left[\left\|h_{g}^{t}\right\|^{2}\right]\leq\frac{4\left(R_{\max}+\delta\right)^{2}}{\lambda_{g}^{2}}.$ (33)
Taking the full expectation of (C.4) and substituting it with (C.4), (33) into (C.4) yields
$\mathbb{E}\left[\left\|\omega^{*}(\theta^{t})-\omega^{t+1}\right\|^{2}\right]$ $\leq\left(1-\frac{\beta\lambda_{g}}{2}\right)\mathbb{E}\left[\left\|\omega^{*}(\theta^{t-1})-\omega^{t}\right\|^{2}\right]+\frac{8\alpha^{2}}{\beta\lambda_{g}^{3}}$
$\quad+\frac{(32\beta^{2}T_{\alpha}+8\alpha\beta+4\beta^{2})(R_{\max}+\delta)^{2}}{\lambda_{g}^{2}}$
$+\frac{16\alpha\beta T_{\alpha}(R_{\max}+\delta)(1+\lambda_{g})}{\lambda_{g}^{2}}.$ (34)
Summing it from $t=T_{\alpha}+1$ to $t=T$ and rearranging the it gives
$\frac{1}{T}\sum_{t=T_{\alpha}}^{T-1}\mathbb{E}\left\|\omega^{*}(\theta^{t})-\omega^{t+1}\right\|^{2}\leq$ $\frac{2}{\beta\lambda_{g}T}\mathbb{E}\left\|\omega^{*}(\theta^{T_{\alpha}})-\omega^{T_{\alpha}+1}\right\|^{2}+\frac{16\alpha^{2}}{\beta^{2}\lambda_{g}^{4}}$
$+\frac{(64\beta T_{\alpha}+16\alpha+8\beta)(R_{\max}+\delta)^{2}}{\lambda_{g}^{3}}$
$+\frac{32\alpha T_{\alpha}(R_{\max}+\delta)(1+\lambda_{g})}{\lambda_{g}^{3}}$
$=$ $\mathcal{O}\left(\frac{1}{\beta T}+\frac{\alpha^{2}}{\beta^{2}}+\beta T_{\alpha}+\alpha+\beta+\alpha T_{\alpha}\right).$ (35)